Evaluate the ∫dx/5-3x


1
Expert's answer
2021-02-16T11:27:28-0500

Solution.

Make a replacement:

"t=5-3x," from here "3x=5-t,
x=-\frac{1}{3}t+\frac{5}{3}."

Then "dx=d(-\frac{1}{3}t+\frac{5}{3})=-\frac{1}{3}dt."

We will have

"\int{\frac{-\frac{1}{3}dt}{t}}=
-\frac{1}{3}\int{\frac{dt}{t}}= 
-\frac{1}{3}\ln{|t|}+C=
-\frac{1}{3}\ln{|{5-3x}|}+C."

Answer.

"\int{\frac{dx}{5-3x}}=
-\frac{1}{3}\ln{|5-3x|}+C"